A 10-foot ladder reaches a height of 8 feet when placed against the side of a building. How far from the base of the building is the ladder?

Answer by checkley77(12844)

a^2+8^2=10^2
a^2+64=100
a^2=100-64
a^2=36
a=sqrt36
a=6 ft. is the distance between the building & the ladder.